# Refining the Complexity Landscape of Speed Scaling: Hardness and Algorithms

### v1

@@ C.2 A Combinatorial Algorithm for FE-IDUU
In this section, we prove Theorem 4.7. Consider an instance of FE-IDUU with jobs $1,\dots,n$ and speeds $s_{1},\dots,s_{k}$. We assume that $r_{1}\leq\dots\leq r_{n}$. Since we have unit-size and unit-weight jobs, ordering the jobs from earliest to latest release time (FIFO) gives an optimal completion ordering. This follows from the fact that the jobs have the same size, and thus processing the jobs in FIFO order means the job being processed always has the least volume left to process, which adheres to the SRPT rule. Hence, the problem is equivalent to finding optimal speeds for completion ordering $1\prec\dots\prec n$. We claim that initiating every job at speed $s_{1}$, and then running the following algorithm, computes an optimal schedule.
See 1
See 4.7
For two speeds, Algorithm 1 is not identical to Algorithm 2. Hence, we would not consider it a direct generalization. However, it is important to note that Barcelo et al. did introduce concepts similar to the affection and shrinking energy: two important concepts for our algorithm. For the runtime of the algorithm, note that a step where we speed up job $j$ ends when either $\kappa_{j}$ or $\Delta_{j}$ changes. If $\kappa_{j}$ changes, then $j$ loses its affection on some job. Thus, this cannot happen more than $n$ times for each job. If $\Delta_{j}$ changes, then $j$ must have reached the next given speed. Hence, this can happen at most $k$ times for each job. It follows that the number of steps is $O(n^{2}k)$. Since we can also compute how much we can increase the speed of a job until it either reaches the next fixed speed or loses affection in polynomial time, it follows that Algorithm 1 runs in polynomial time.
To prove the correctness of Algorithm 1, we will show that for any schedule that was not constructed by the algorithm we can do the following. By slowing down one job and speeding up another, we either obtain an improvement in the objective, or a schedule that is “more similar" to the schedule constructed by Algorithm 1. To show that such an exchange exists, we exploit the fact that affection is easier to describe in a setting with unit-size and unweighted jobs. To use this, we introduce new terminology and prove several results on the affection in this setting. We then define a way to construct any structure that is similar to Algorithm 1. This will allow us to compare schedules to the one from the algorithm, and determine which jobs we need for the exchange.
Since the jobs are ordered from earliest to latest release time, jobs are never preempted. Hence, the affection set of a job $j$ is always of the form $\{j,j+1,\dots,j+\ell\}$ where we have $\ell\in\{0,\dots,n-1\}$. The following definition captures this and will help us describe affection throughout the algorithm.
@@
For a schedule $\sigma$, an affection chain is a maximal sequence of jobs $(j,j+1,\dots,j+\ell)$ such that for all $i\in\{0,\dots,\ell-1\}$, job $j+i$ affects job $j+i+1$.
@@
Let $\sigma$ and $\sigma^{\prime}$ be two different schedules. For jobs $j_{1},j_{2}$, if $\mathcal{K}_{j_{1}}(\sigma)=\mathcal{K}_{j_{2}}(\sigma)$ and $\mathcal{K}_{j_{1}}(\sigma^{\prime})=\mathcal{K}_{j_{2}}(\sigma^{\prime})$, then
$\kappa_{j_{1}}(\sigma)-\kappa_{j_{2}}(\sigma)=\kappa_{j_{1}}(\sigma^{\prime})-\kappa_{j_{2}}(\sigma^{\prime})=j_{2}-j_{1}\,.$
If $\mathcal{K}^{+}_{j_{1}}(\sigma)=\mathcal{K}^{+}_{j_{2}}(\sigma)$ and $\mathcal{K}^{+}_{j_{1}}(\sigma^{\prime})=\mathcal{K}^{+}_{j_{2}}(\sigma^{\prime})$, then
$\kappa^{+}_{j_{1}}(\sigma)-\kappa^{+}_{j_{2}}(\sigma)=\kappa^{+}_{j_{1}}(\sigma^{\prime})-\kappa^{+}_{j_{2}}(\sigma^{\prime})=j_{2}-j_{1}\,.$
@@
Let $\sigma$ and $\sigma^{\prime}$ be two different schedules and let $j_{1}<j_{2}$. If $\mathcal{K}_{j_{1}}(\sigma)=\mathcal{K}_{j_{2}}(\sigma)$, then
@@
Consider some schedule $\sigma$ and jobs $j_{1},j_{2}$. For an $\varepsilon>0$, let $\sigma^{+\varepsilon}_{j_{1}}$ be a schedule with $x_{j}(\sigma^{+\varepsilon}_{j_{1}})=x_{j}(\sigma)$ for all $j\neq j_{1}$ and $x_{j_{1}}(\sigma^{+\varepsilon}_{j_{1}})=x_{j_{1}}(\sigma)+\varepsilon$. Let $\sigma^{-\varepsilon}_{j_{2}}$ be a schedule with $x_{j}(\sigma^{-\varepsilon}_{j_{2}})=x_{j}(\sigma)$ for all $j\neq j_{2}$ and $x_{j_{2}}(\sigma^{-\varepsilon}_{j_{2}})=x_{j_{2}}(\sigma)-\varepsilon$. Finally, let $\sigma_{j_{1}j_{2}}^{\varepsilon}$ be the schedule with
@@
Let $j+\ell$ be the last job in $\mathcal{K}^{+}_{j}(\sigma)$. Recall that by the definition of lower affection, $C_{j+i}(\sigma)\geq r_{j+i+1}$ for all $i\in\{0,\dots,\ell-1\}$. Furthermore, have that $C_{j+\ell}(\sigma)<r_{j+\ell+1}$. Hence, if we choose
@@
then $C_{j+i}(\sigma_{j}^{+\varepsilon})\geq r_{j+i+1}+\varepsilon>r_{j+i+1}$ for all $i\in\{0,\dots,\ell-1\}$, and $C_{j+\ell}(\sigma_{j}^{+\varepsilon})\leq r_{j+\ell+1}$. Thus, $\kappa_{j}(\sigma_{j}^{+\varepsilon})=\kappa_{j}^{+}(\sigma)$. Note that if $j+\ell=n$, then the statement holds for any $\varepsilon>0$.
In $\sigma^{+\varepsilon}_{j_{1}}$, the affection of job $j_{2}$ might also change compared to $\sigma$. Hence, describing how the objective changes from $\sigma$ to $\sigma_{j_{1}j_{2}}^{\varepsilon}$ is not always straightforward. The following lemma will help us identify when $\sigma_{j_{1}j_{2}}^{\varepsilon}$ gives an improvement. For a schedule $\sigma$, let $\mathrm{Obj}(\sigma)$ denote the objective value of $\sigma$.
@@
Consider some schedule $\sigma$ and jobs $j_{1},j_{2}$. Suppose one of the following holds:
@@
  • $j_{1}<j_{2}$ and $\mathcal{K}^{+}_{j_{1}}(\sigma)=\mathcal{K}^{+}_{j_{2}}(\sigma)$.
@@
$\mathrm{Obj}(\sigma^{\varepsilon}_{j_{1}j_{2}})\leq\mathrm{Obj}(\sigma)-\varepsilon\left(\left(\kappa^{+}_{j_{2}}(\sigma)-\Delta_{j_{2}}(\sigma)\right)-\left(\kappa^{+}_{j_{1}}(\sigma)-\Delta^{+}_{j_{1}}(\sigma)\right)\right)\,.$
@@
Suppose at least one of the two given requirements holds. If $\mathcal{K}_{j_{1}}(\sigma)=\mathcal{K}_{j_{2}}(\sigma)$, then $j_{1}$ and $j_{2}$ remain in the same affection chain in $\sigma_{j_{1}}^{+\varepsilon}$ for any $\varepsilon>0$. If $j_{1}<j_{2}$ and $\mathcal{K}^{+}_{j_{1}}(\sigma)=\mathcal{K}^{+}_{j_{2}}(\sigma)$, then by Observation C.2 we can chose an $\varepsilon^{\prime}>0$ small enough such that $\kappa_{j_{1}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})=\kappa_{j_{1}}^{+}(\sigma)$. Hence, we also obtain that $\mathcal{K}_{j_{1}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})=\mathcal{K}^{+}_{j_{1}}(\sigma)$. Thus, we have that for both cases $\mathcal{K}_{j_{1}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})=\mathcal{K}_{j_{2}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})$. By Observation C.2, and using that $\kappa_{j_{1}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})=\kappa_{j_{1}}^{+}(\sigma)$, we obtain that
@@
$\mathrm{Obj}(\sigma^{\varepsilon}_{j_{1}j_{2}})$ $=\mathrm{Obj}(\sigma)+\varepsilon\left(\kappa_{j_{1}}^{+}(\sigma)-\Delta_{j_{1}}^{+}(\sigma)\right)-\varepsilon\left(\kappa_{j_{2}}(\sigma_{j_{1}}^{\varepsilon^{\prime}})-\Delta_{j_{2}}(\sigma_{j_{1}}^{\varepsilon^{\prime}})\right)$
@@
Note that any schedule $\sigma$ can be constructed by initiating every job at speed $s_{1}$, and then speeding up the jobs until every job $j$ runs at speed $s_{j}(\sigma)$. We could do these speed-ups in many different orders, but to better compare a schedule to the schedule obtained by Algorithm 1, we define a way to construct a schedule that is similar to the algorithm.
A construction of a schedule $\sigma$ consists of several steps $(\sigma^{0},\dots,\sigma^{T})$ that satisfy the following
@@
    For all $\tau\in\{1,\dots,T\}$, there is exactly one $j^{\tau}\in\{1,\dots,n\}$ with $s_{j^{\tau}}(\sigma^{\tau-1})<s_{j^{\tau}}(\sigma^{\tau})$. For all $j\neq j^{\tau}$, we have that $s_{j}(\sigma^{\tau-1})=s_{j}(\sigma^{\tau})$.
@@
Since Algorithm 1 always speeds up a job with maximum $\kappa-\Delta$, its steps can also be considered a construction. Generally,in a construction, we are allowed to speed up a job which does not have maximum $\kappa-\Delta$, but only if all jobs with maximum $\kappa-\Delta$ already run at their final speed. If we do speed up such a job, then the construction can be the same as the construction from the Algorithm. Hence, we are interested in comparing constructions that do speed a job with non-maximum $\kappa-\Delta$, to constructions that never speed up such a job. For this purpose, we introduce the following:
@@
By Corollary 2.10, a construction of an optimal schedule cannot end after step $\tau$ if there is a job $j$ with $\kappa_{j}(\sigma^{\tau})-\Delta_{j}>0(\tau)$. However, it is not clear whether we should speed up jobs with $\kappa_{j}-\Delta_{j}=0$. To better compare different constructions, we may assume the following:
@@
To prove the correctness of Algorithm 1, we first prove the following: If we have a schedule with a construction that does not follow the $(\kappa-\Delta)$-rule, then we can find two jobs, where slowing down one job and speeding up the other leads to an improvement in the objective value. Hence, such a schedule can never be optimal. To be more specific: the job we speed up will be one that would have adhered to $(\kappa-\Delta)$-rule in a construction step where a violation occurred. The main challenge lies in identifying the two jobs for this exchange, and quantifying the change in the objective. Hence, we introduce the following lemma.
@@
for all $j^{\prime}\in\mathcal{K}_{j}(\sigma^{\tau-1})$. Furthermore suppose $\kappa_{j}(\sigma^{\tau-1})-\Delta_{j}(\sigma^{\tau-1})>0$, $\kappa_{j}(\sigma^{\tau})-\Delta_{j}(\sigma^{\tau})\leq 0$, and $j^{\tau}\neq j$. If all steps after $\tau$ follow the $(\kappa-\Delta)$-rule, then there is an $\varepsilon>0$ such that
@@
$\kappa_{j}(\sigma^{\tau^{\prime}-1})-\Delta_{j}(\sigma^{\tau^{\prime}-1})\geq\kappa_{j^{\prime}}(\sigma^{\tau^{\prime}-1})-\Delta_{j^{\prime}}(\sigma^{\tau^{\prime}-1})$
@@
$\kappa_{j}(\sigma^{\tau^{\prime}-1})-\Delta_{j}(\sigma^{\tau^{\prime}-1})>\kappa_{j^{\tau^{\prime}}}(\sigma^{\tau^{\prime}-1})-\Delta_{j^{\tau^{\prime}}}(\sigma^{\tau^{\prime}-1})>0$
@@
From Lemma C.11, we obtain that there is an $\varepsilon>0$ such that $\mathrm{Obj}(\sigma_{j^{\tau}j})^{\varepsilon}<\mathrm{Obj}(\sigma)$, which contradicts that $\sigma$ is optimal.
While Corollary C.13 shows the construction of an optimal schedule must follow the $(\kappa-\Delta)$-rule, due to tiebreakers, there can still be many different schedules with constructions that do follow this rule. Thus, we cannot yet conclude that the schedule constructed by Algorithm 1 is indeed optimal. to complete our proof, we show that any schedule with a construction that follows the $(\kappa-\Delta)$-rule that also satisfies the conditions from Observation C.2 is optimal. To do this, we use Lemma C.11 to show that for two schedules that both satisfy these conditions, we can make one of them “more similar" to the other, without changing its objective value. Hence, their objective value must be the same. To describe more concretely what “more similar" means, we introduce the following notation.
@@
Since Algorithm 1 always follows the $(\kappa-\Delta)$-rule and yields a schedule $\sigma$ that satisfies $\kappa_{j}(\sigma)-\Delta_{j}(\sigma)\leq 0$ and $\kappa^{+}_{j}(\sigma)-\Delta^{+}_{j}(\sigma)>0$ for all $j\in\{1,\dots,n\}$, Theorem 4.7 follows directly from Lemma C.15.

### v2

@@ C.2 A Combinatorial Algorithm for FE-IDUU
In this section, we prove Theorem 4. Recall that, since a FIFO ordering is always optimal, we use only job speeds to describe a schedule. We claim that initializing every job at speed $s_{1}$, and then running the following algorithm computes an optimal schedule.
while There is a job $j$ with $\kappa_{j}\geq\Delta_{j}$ do
Algorithm 4 The $(\kappa-\Delta)$-Algorithm.
See 5.1
To prove the correctness of Algorithm 4, we show that for any schedule that was not constructed by the algorithm, we can do the following. By slowing down one job and speeding up another, we either obtain an improvement in the objective, or a schedule that is “more similar” to the schedule constructed by Algorithm 4. To show that such an exchange exists, we exploit the fact that affection is easier to describe in a setting with unit-size and unweighted jobs. To use this, we introduce new terminology and prove several results on the affection in this setting. Algorithm 4 starts with all jobs at speed $s_{1}$ and then increases the speed of the jobs in a specific order. To better compare schedules to the one from the algorithm, we define a way to construct schedules that is similar to the algorithm, but can be used to reconstruct any given schedule. We use this construction to determine which jobs we need for the exchange.
Since the jobs are unit-size and are ordered from earliest to latest release time, jobs are never preempted. Hence, the affection set of a job $j$ is always of the form $\{j,j+1,\dots,j+\ell\}$ where we have $\ell\in\{0,\dots,n-1\}$. The following definition captures this and will help us describe affection throughout the algorithm.
@@
For a schedule $\sigma$, an affection chain is a maximal sequence of jobs $(j,\dots,j+\ell)$ such that job $j$ affects job $j+\ell$.
@@
For a schedule $\sigma$ and jobs $j_{1},j_{2}$, if $\mathcal{K}_{j_{1}}(\sigma)=\mathcal{K}_{j_{2}}(\sigma)$, then
$\kappa_{j_{1}}(\sigma)-\kappa_{j_{2}}(\sigma)=j_{2}-j_{1}\,.$
If $\mathcal{K}^{+}_{j_{1}}(\sigma)=\mathcal{K}^{+}_{j_{2}}(\sigma)$, then
$\kappa^{+}_{j_{1}}(\sigma)-\kappa^{+}_{j_{2}}(\sigma)=j_{2}-j_{1}\,.$
@@
For schedules $\sigma$ and $\sigma^{\prime}$ and jobs $j_{1}<j_{2}$, if $\mathcal{K}_{j_{1}}(\sigma)=\mathcal{K}_{j_{2}}(\sigma)$, then
@@
Consider some schedule $\sigma$ and jobs $j_{1},j_{2}$. For any $\varepsilon>0$, let $\sigma^{+\varepsilon}_{j_{1}}$ be a schedule with $x_{j}(\sigma^{+\varepsilon}_{j_{1}})=x_{j}(\sigma)$ for all $j\neq j_{1}$ and $x_{j_{1}}(\sigma^{+\varepsilon}_{j_{1}})=x_{j_{1}}(\sigma)+\varepsilon$. Let $\sigma^{-\varepsilon}_{j_{2}}$ be a schedule with $x_{j}(\sigma^{-\varepsilon}_{j_{2}})=x_{j}(\sigma)$ for all $j\neq j_{2}$ and $x_{j_{2}}(\sigma^{-\varepsilon}_{j_{2}})=x_{j_{2}}(\sigma)-\varepsilon$. Finally, let $\sigma_{j_{1}j_{2}}^{\varepsilon}$ be the schedule with
@@
Let $j+\ell$ be the last job in $\mathcal{K}^{+}_{j}(\sigma)$, and thus $\kappa_{j}^{+}(\sigma)=\ell+1$. If $j+\ell=n$, then the statement holds for any $\varepsilon>0$. Otherwise, by the definition of lower affection, we have $C_{j+i}(\sigma)\geq r_{j+i+1}$ for all $i\in\{0,\dots,\ell-1\}$ and $C_{j+\ell}(\sigma)<r_{j+\ell+1}$. Let
@@
then $C_{j+i}(\sigma_{j}^{+\varepsilon})=r_{j+\ell+1}\geq r_{j+i+1}+\varepsilon>r_{j+i+1}$ for all $i\in\{0,\dots,\ell-1\}$, and $C_{j+\ell}(\sigma_{j}^{+\varepsilon})\leq r_{j+\ell+1}$. Thus, $\kappa_{j}(\sigma_{j}^{+\varepsilon})=\ell+1=\kappa_{j}^{+}(\sigma)$.
The affection of job $j_{2}$ in $\sigma^{+\varepsilon}_{j_{1}}$ may differ compared to $\sigma$. Hence, describing how the objective changes from $\sigma$ to $\sigma_{j_{1}j_{2}}^{\varepsilon}$ is not always straightforward. The following lemma helps us identify when $\sigma_{j_{1}j_{2}}^{\varepsilon}$ gives an improvement. For a schedule $\sigma$, let $\mathrm{Obj}(\sigma)$ denote the objective value of $\sigma$.
@@
Consider some schedule $\sigma$ and jobs $j_{1},j_{2}$. If at least one of the following conditions holds:
@@
  • $j_{1}<j_{2}$ and $\mathcal{K}^{+}_{j_{1}}(\sigma)=\mathcal{K}^{+}_{j_{2}}(\sigma)$,
@@
$\mathrm{Obj}(\sigma^{\varepsilon}_{j_{1}j_{2}})=\mathrm{Obj}(\sigma)-\varepsilon\left(\left(\kappa^{+}_{j_{2}}(\sigma)-\Delta_{j_{2}}(\sigma)\right)-\left(\kappa^{+}_{j_{1}}(\sigma)-\Delta^{+}_{j_{1}}(\sigma)\right)\right)\,.$
@@
First, note that if $\mathcal{K}_{j_{1}}(\sigma)=\mathcal{K}_{j_{2}}(\sigma)$, then $j_{1}$ and $j_{2}$ remain in the same affection chain in $\sigma_{j_{1}}^{+\varepsilon}$ for any $\varepsilon>0$. Next, assume that $j_{1}<j_{2}$ and $\mathcal{K}^{+}_{j_{1}}(\sigma)=\mathcal{K}^{+}_{j_{2}}(\sigma)$. By Observation C.2 we can chose an $\varepsilon^{\prime}>0$ small enough such that $\kappa_{j_{1}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})=\kappa_{j_{1}}^{+}(\sigma)$. Since $j_{1}<j_{2}$, it follows that $j_{2}\in\mathcal{K}_{j_{1}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})$.
Thus, if one of the two conditions holds, we have that $\mathcal{K}_{j_{1}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})=\mathcal{K}_{j_{2}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})$. By Observation C.2, and using that $\kappa_{j_{1}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})=\kappa_{j_{1}}^{+}(\sigma)$, we obtain that
@@
$\mathrm{Obj}(\sigma^{\varepsilon}_{j_{1}j_{2}})$ $=\mathrm{Obj}(\sigma)+\varepsilon\left(\kappa_{j_{1}}^{+}(\sigma)-\Delta_{j_{1}}^{+}(\sigma)\right)-\varepsilon\left(\kappa_{j_{2}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})-\Delta_{j_{2}}(\sigma_{j_{1}}^{+\varepsilon^{\prime}})\right)$
@@
Note that any schedule $\sigma$ can be constructed by initializing every job at speed $s_{1}$, and then speeding up the jobs until every job $j$ runs at speed $s_{j}(\sigma)$. We could do these speed-ups in many different orders, but to better compare a schedule to the schedule obtained by Algorithm 4, we define a way to construct a schedule that speeds up the jobs in an order similar to the algorithm.
A construction of a schedule $\sigma$ consists of several steps $1,\dots,\tau$ with corresponding intermediate schedules $(\sigma^{0},\dots,\sigma^{T})$ that satisfy the following
@@
    For all $\tau\in\{1,\dots,T\}$, there is exactly one $j^{\tau}\in\{1,\dots,n\}$ with $s_{j^{\tau}}(\sigma^{\tau-1})<s_{j^{\tau}}(\sigma^{\tau})$. For all $j\neq j^{\tau}$, we have that $s_{j}(\sigma^{\tau-1})=s_{j}(\sigma^{\tau})$. We say that we speed up $j^{\tau}$ in step $\tau$ of the construction.
@@
Since Algorithm 4 always speeds up a job with maximum $\kappa-\Delta$, its steps also constitute a construction. Generally, in a construction, we are allowed to speed up a job which does not have maximum $\kappa-\Delta$, but only if all jobs with maximum $\kappa-\Delta$ already run at their final speed. If we do speed up such a job, then the construction can be the same as the construction from the Algorithm. Hence, we are interested in comparing constructions that do speed up a job with non-maximum $\kappa-\Delta$, to constructions that never speed up such a job. For this purpose, we introduce the following:
@@
By Corollary 2.10, a construction of an optimal schedule cannot end after step $\tau$ if there is a job $j$ with $\kappa_{j}(\sigma^{\tau})-\Delta_{j}(\sigma^{\tau})>0$. However, it is not clear whether we should speed up jobs with $\kappa_{j}-\Delta_{j}=0$. To better compare different constructions, we may assume the following:
@@
To prove the correctness of Algorithm 4, we first prove the following: If we have a schedule with a construction that does not follow the $(\kappa-\Delta)$-rule, then we can find two jobs where slowing down one job and speeding up the other leads to an improvement in the objective value. Thus, such a schedule can never be optimal. To be more specific: the job we speed up will be one that would have adhered to the $(\kappa-\Delta)$-rule in a construction step where a violation occurred. The main challenge lies in identifying the job that should be slowed down in this exchange. Hence, we introduce the following lemma.
@@
$\kappa_{j}(\sigma^{\tau-1})-\Delta_{j}(\sigma^{\tau-1})\geq\kappa_{j^{\prime}}(\sigma^{\tau-1})-\Delta_{j^{\prime}}(\sigma^{\tau-1})\,$,
for all $j^{\prime}\in\mathcal{K}_{j}(\sigma^{\tau-1})$. Furthermore, suppose $\kappa_{j}(\sigma^{\tau-1})-\Delta_{j}(\sigma^{\tau-1})>0$, $\kappa_{j}(\sigma^{\tau})-\Delta_{j}(\sigma^{\tau})\leq 0$, and $j^{\tau}\neq j$. If all steps after $\tau$ follow the $(\kappa-\Delta)$-rule, then there is an $\varepsilon>0$ such that
@@
$\kappa_{j}(\sigma^{\tau^{\prime}-1})-\Delta_{j}(\sigma^{\tau^{\prime}-1})\geq\kappa_{j^{\prime}}(\sigma^{\tau^{\prime}-1})-\Delta_{j^{\prime}}(\sigma^{\tau^{\prime}-1})\,$,
@@
$\kappa_{j}(\sigma^{\tau^{\prime}-1})-\Delta_{j}(\sigma^{\tau^{\prime}-1})>\kappa_{j^{\tau^{\prime}}}(\sigma^{\tau^{\prime}-1})-\Delta_{j^{\tau^{\prime}}}(\sigma^{\tau^{\prime}-1})>0\,$,
@@
From Lemma C.11, we obtain that there is an $\varepsilon>0$ such that $\mathrm{Obj}(\sigma_{j^{\tau}j})<\mathrm{Obj}(\sigma)$, which contradicts that $\sigma$ is optimal.
While Corollary C.13 shows it is necessary for a construction of an optimal schedule to follow the $(\kappa-\Delta)$-rule, it remains to show that it is sufficient (if the schedule also satisfies the conditions from Observation C.2)). Thus, we cannot yet conclude that the schedule constructed by Algorithm 4 is indeed optimal. To complete our proof, we show that any schedule with a construction that follows the $(\kappa-\Delta)$-rule that also satisfies the conditions from Observation C.2 is optimal. To do this, we use Lemma C.11 to show that for two schedules that both satisfy these conditions, we can make one of them “more similar" to the other, without changing its objective value. Hence, their objective value must be the same. To describe more concretely what “more similar" means, we introduce the following notation.
@@
Since Algorithm 4 always follows the $(\kappa-\Delta)$-rule and yields a schedule $\sigma$ that satisfies $\kappa_{j}(\sigma)-\Delta_{j}(\sigma)\leq 0$ and $\kappa^{+}_{j}(\sigma)-\Delta^{+}_{j}(\sigma)>0$ for all $j\in\{1,\dots,n\}$, the correctness of Algorithm 4 follows directly from Lemma C.15.
Finally, let us consider the running time of Algorithm 4. Since the steps of the algorithm constitute a construction, it follows from Observation C.2 for any job $j$ we have that $\kappa_{j}$ never increases and $\Delta_{j}$ never decreases. Thus, a step where we speed up job $j$ ends when either $\kappa_{j}$ or $\Delta_{j}$ decreases. Since the affection of any job $j$ is upper bounded by $n$, it follows that $\kappa_{j}$ cannot decrease more than $n$. Similarly, $\Delta_{j}$ cannot increase more than $k$. It follows that the number of steps is $O(n^{3}k)$. Since we can also compute in polynomial time how much we can increase the speed of a job until it either reaches the next fixed speed or loses affection, it follows that Algorithm 4 runs in polynomial time, proving Theorem 5.1.